24. While observing the obelisk in Saint Peter’s Square in Vatican City, you stand at the edge of the shadow it is casting across the ground, which is about 120 m long. While kneeling, you estimate your line of sight with the top of the obelisk to be about 12o above the horizontal. Approximately how tall is the obelisk? (3)

A) 9.6 m C) 23.1 m E) 26 m 

B) 120 m D) 190 m 

25. A partly full paint can has 2.75 liters of paint left in it. What is volume of the paint in cubic centimeters? (3)

A) 2.75 x 10-3 cm3 C) 275 x 10-3 cm3 E) 27.5 x 103 cm3

B) 2.75 cm3 D) 2.75 x 103















1
Expert's answer
2021-06-12T10:42:38-0400

24) We can find the height of the obelisk from the geometry:


"tan\theta=\dfrac{h}{L},""h=Ltan\theta=120\ m\cdot tan12^{\circ}=26\ m."

Therefore, the correct answer is (E) 26 m.

25)


"2.75\ L\cdot\dfrac{0.001\ m^3}{1\ L}\cdot\dfrac{10^6\ cm^3}{1\ m^3}=2750\ cm^3=2.75\cdot10^3\ cm^3."

Therefore, the correct answer is (D) 2.75 x 103 cm3